# Enhancing Automated Loop Invariant Generation for Complex Programs with Large Language Models

### v1

@@ 4 Experiment
LIG-MM SVCOMP24 FFmpeg Frama-C SyGuS OOPSLA SVCOMP22 Struc non-Struc All
Traditional tools
@@
4.1 Experimental Setup
Dataset Our dataset is mainly divided into two parts, those with data structures and those without data structures. The dataset with data structure consists of three parts: LIG-MM Liu et al. (2024); Thinklab-SJTU (2023), SV-COMP Beyer (2024); SoSy-Lab (2024), and FFmpeg FFmpeg (2023). We will elaborate on these three datasets in more detail later. The dataset without data structure mainly comes from the benchmark from AutoSpec Wen et al. (2024), and we will also briefly describe this dataset.
First, we selected the dataset from LIG-MM, which is a new benchmark specifically for programs with complex data structures and memory manipulations. LIG-MM is also divided into four parts: (1) the course programs; (2) SV-COMP22 Beyer (2022); (3) Benchmark from SLING Le et al. (2019), which is a separation logic loop invariant inferring tool using dynamic analysis; (4) Some real-world programs from well-known software and systems. LIG-MM consists of C code and focuses on some complex data structures, especially the single-linked list, double-linked list, tree, and hash table.
We also picked C programs from SV-COMP 2024. LIG-MM and the benchmark of AutoSpec both select some programs from SV-COMP. Unlike them, we selected 31 programs, all of which include at least one loop and one data structure. Among these data structures, we excluded all structures containing the single-linked list, double-linked list, tree, and hash table to avoid duplication with the selected parts in LIG-MM.
To evaluate the effectiveness of ACInv on real-world programs, we selected several program segments from the widely used audio and video processing library FFmpeg. Given the complexity of function calls and structures within C programs, as well as the extensive codebase, we focused on specific portions of the program. We applied simple partitioning techniques, ensuring that ACInv could concentrate on the relevant functions for analysis, and we picked the programs with at least one loop and at least one data structure.
We also selected programs from the benchmark of AutoSpec, which assesses the tool’s ability to synthesize specifications, including loop invariants, preconditions, and postconditions of functions. Notably, some C programs in this benchmark do not contain loops, so we specifically chose a subset that includes loops for our testing. The benchmark from AutoSpec comprises four parts: Frama-C, SyGuS, OOPSLA-13, and SV-COMP22. All of these components consist of C code and are primarily focused on numerical programs without explicit structural declarations.
@@ 4.1 Experimental Setup
Baseline In this study, we compare the performance of several tools with ACInv on datasets both with and without data structures. For datasets with data structures, we evaluate SLING, LLM-SE Liu et al. (2024), GPT-4o, and AutoSpec. SLING is a dynamic analysis tool for inferring loop invariants based on separation logic. LLM-SE leverages a fine-tuned language model (LLM) to generate separation logic loop invariants, incorporating predicate decomposition into its training. Since it neither provides the fine-tuned model nor the training data, we only use the results of its open-source data as a reference. AutoSpec employs hierarchical specification generation to produce ACSL-style annotations. Both LLM-SE and AutoSpec utilize verification checkers to ensure the correctness of the generated invariants. GPT-4o, in this context, is used by directly embedding the code into prompts, asking GPT-4o-2024-08-06 to generate loop invariants without employing specific prompting techniques or examples.
@@
For datasets without data structures, we assess CODE2Inv Si et al. (2018, 2020), CLN2Inv Ryan et al. (2020), GPT-4o, and AutoSpec. CODE2Inv and CLN2Inv are state-of-the-art invariant generation tools for numerical programs, based on graph neural networks and continuous logic networks, respectively. AutoSpec was originally a tool for testing on this dataset, and we naturally referred to its results.
4.2 Overview of results
Table 1 shows the main results of our experiments, where completion means the rate of programs whose loop invariants can be inferred correctly to complete the verification of this program. In this experiment, considering the balance between time and performance, we set the augmentation iteration parameter $k$ to 5.
@@ 4.2 Overview of results
We first focus on datasets that include data structures. Compared to the traditional tool SLING, LLM-based technologies, including ACInv, demonstrate strong applicability. While SLING can only reason about the linked list portion of LIG-MM, it struggles with other complex data structures. Experimental results from LLM-SE indicate good performance on singly linked lists (sll), doubly linked lists (dll), trees, and hash tables. However, LLM-SE did not provide training sets or fine-tuned models, limiting the ability to evaluate its performance on other data structures. Additionally, LLM-SE incorporates pre-defined predicates during fine-tuning, enabling abstraction during reasoning, though this approach may restrict its effectiveness on other data structures or non-pointer data structures.
@@
In comparison to pure LLM querying and the latest technology, AutoSpec, ACInv performs well on datasets involving data structures. This performance is anticipated. In pure LLM querying, we observed that while the LLM can recognize data structures in the program and analyze the associated variables, it struggles to abstract them into suitable logical forms. This often results in accurate formulas being expressed in natural language rather than logical terms. Regarding AutoSpec, its examples primarily focus on numerical types, leading to suboptimal performance on datasets involving data structures.
@@
4.3 Ablation Study
@@
We conducted an ablation study on the final augmentation stage. Figure 4 shows the results of the experiment on the augmentation of ACInv. We tested the invariants generated by ACInv on a structured dataset, where Acc means the proportion of generated invariants that can pass the Frama-c check, and model@k means $k$ augmentation iterations were performed on the model. The experiment shows that when we use appropriate prompts and examples, the augmentation stage can judge the correctness of the invariants with a relatively high probability and filter out incorrect invariants.
@@ 5 Conclusion
6 Limitations
Appendix A Examples of generating predicates
struct stack_int{
size_t top;
int data[MAX_SIZE];
(a) An non-recursive data structure example: stack
struct Node{
int data;
struct Node * next;
(b) A recursive data structure example: linked list
Figure 5: Two examples of data structure.
The C language definition of the stack and linked list is usually as Figure 5.
[…]
The recursive data structure is usually related to pointers. Taking inspiration from separation logic Reynolds (2002), we provide such a recursive definition for the following single linked list, which considered both empty and non-empty linked lists, and expressed that the pointer p1 to p2 is a single linked list such as Figure 7.
predicate
valid_stack_int(struct stack_int* s) =
[…]
predicate
full_stack_int(struct stack_int* s) =
[…]
predicate
empty_stack_int(struct stack_int* s) =
[…]
Figure 6: Predicates of stack
inductive is_linked_list{L}
(struct Node *p1, struct Node *p2) {
case empty{L}:
$\forall$ struct Node *p1, *p2;
(p1 == p2) ==> is_linked_list(p1, p2);
case non_empty{L}:
$\forall$ struct Node *p1, *p2;
[…]
is_linked_list(p1->next, p2)
) ==> is_linked_list(p1, p2);
Figure 7: Predicates of linked list

### v3

@@ 4.2 Experimental Setup
LIG-$M$ SVCOMP24 FFmpeg Frama-C SyGuS OOPSLA SVCOMP22 Struc non-Struc All
non-LLM
@@
Baseline In this study, we compare the performance of several tools with ACInv on datasets both with and without data structures. For datasets with data structures, we evaluate SLING, LLM-SE [23], GPT-4o, and AutoSpec. SLING is a dynamic analysis tool for inferring loop invariants based on separation logic. LLM-SE leverages a fine-tuned LLM (which is CodeGen [43] in the specific implementation) to generate separation logic loop invariants, incorporating predicate decomposition into its training. AutoSpec employs hierarchical specification generation to produce ACSL-style annotations. Both LLM-SE and AutoSpec utilize verification checkers to ensure the correctness of the generated invariants. GPT-4o, in this context, is used by directly embedding the code into prompts, asking GPT-4o-2024-08-06 to generate loop invariants without employing specific prompting techniques or examples.
@@
For datasets without data structures, we assess CODE2Inv [16, 44], CLN2Inv [14], GPT-4o, and AutoSpec. CODE2Inv and CLN2Inv are state-of-the-art invariant generation tools for numerical programs, based on graph neural networks and continuous logic networks, respectively. AutoSpec was originally a tool for testing on this benchmark, and we naturally referred to its results.
Environment We primarily employ GPT-4o-2024-08-06 as the core model for implementing ACInv, with GPT-3.5-turbo serving as a control in our experiments. The experiments are performed on a system equipped with an AMD Ryzen 9 5950X CPU, running Ubuntu 23.04, 128GB of RAM, and two NVIDIA RTX 4090 GPUs, each with 24GB of memory.
4.3 RQ1: Performance on Structure Datasets
@@
Table 1 shows the main results of our experiments, where completion means the rate of programs whose loop invariants can be inferred correctly to complete the verification of this program. In this experiment, considering the balance between time and performance, we set the augmentation iteration parameter k to 5. Compared to the traditional tool SLING, LLM-based technologies, including ACInv, demonstrate strong applicability. While SLING can only reason about the linked list portion of LIG-MM, it struggles with other complex data structures. Experimental results from LLM-SE indicate good performance on singly linked lists (sll), doubly linked lists (dll), trees, and hash tables. However, LLM-SE did not provide training sets or fine-tuned models, limiting the ability to evaluate its performance on other data structures. Additionally, LLM-SE incorporates pre-defined predicates during fine-tuning, enabling abstraction during reasoning, though this approach may restrict its effectiveness on other data structures or non-pointer data structures.
@@
In comparison to pure LLM querying and the latest technology, AutoSpec, ACInv performs well on datasets involving data structures. This performance is anticipated. In pure LLM querying, we observed that while the LLM can recognize data structures in the program and analyze the associated variables, it struggles to abstract them into suitable logical forms. This often results in accurate formulas being expressed in natural language rather than logical terms. Regarding AutoSpec, its examples primarily focus on numerical types, leading to suboptimal performance on benchmarks involving data structures.
@@ 4.4 RQ2: Performance on non-Structure Datasets
4.5 Case Study:
Figure 8: In-place reversal of the singly linked list
Considering the example as Figure 8. This program snippet performs an in-place iterative reversal of a singly linked list through pointer manipulation. Three pointers are initialized: old references the head of the original list, newn is set to NULL, representing the nascent reversed list, and curr is initialized to traverse the original list starting at old. During each iteration, old is advanced to its successor node to preserve traversal integrity, the next pointer of curr is redirected to newn to reverse linkage direction, newn is updated to reference the current node, thus extending the reversed sublist, and curr is reset to old for subsequent processing. The loop terminates when curr becomes NULL, at which point newn references the head of the fully reversed list. To give a specific example, if this linked list was originally A->B->C->NULL, it becomes C->B->A->NULL after the reversal.
Most of the other methods based on LLM can only obtain one invariant for this loop, namely old == curr. By observing the detailed output of the LLM, we can find that the LLM can actually capture some more specific properties. That is, this linked list is divided into two parts from the pointers old and curr. The part before curr is the flipped linked list, while the part after old is the linked list that has not been flipped yet. However, LLMs cannot describe this property completely formally. Instead, it describes it in a semi-formal and semi-natural language way, such as curr == suffix of the original list starting at position i. And our work can provide a more formal description of this nature, namely is_linked_list(curr,NULL), is_linked_list(old,NULL).
4.6 RQ3: Testing for the Augmentation
@@
Figure 9 shows the results of the experiment on the augmentation of ACInv. We tested the invariants generated by ACInv on a structured dataset, where Acc means the proportion of generated invariants that can pass the frama-c check, and model@k means k augmentation iterations were performed on the model. The experiment shows that when we use appropriate prompts and examples, the augmentation stage can judge the correctness of the invariants with a relatively high probability and filter out incorrect invariants.
4.7 Threats to validity
5 Related Work
5.1 LLM for static analysis
Recent advances in LLMs have opened new avenues in static analysis. By integrating LLMs with traditional static analysis techniques, their scope can be broadened, leading to more effective and deeper program comprehension.
LLM4SA [45] leverages LLMs to analyze static warnings, reducing the need for manual effort. LLift [18] integrates static analysis with LLM-based prompting, successfully identifying Use Before Initialization (UBI) bugs in the Linux kernel. CFStra [19] employs LLMs to automatically select verification strategies based on code features and specifications. IRIS [46] combines LLMs with static analysis to perform whole-repository reasoning, detecting security vulnerabilities. NS-Slicer [47] uses a fine-tuned model to predict static program slices for both complete and partial code. Codeplan [48] synthesizes a multi-step chain of edits to assist with repository-level coding. ClarifyGPT [49] enhances code generation by leveraging LLMs, while AutoFL [50] utilizes LLMs to generate explanations of bugs for the purpose of Fault Localization (FL). Overall, incorporating LLMs has mitigated challenges related to applicability and performance overhead in static analysis and has enhanced the ability to analyze real-world code.
5.2 Invariants Generation
Symbolic methods Traditional symbolic methods for invariant generation typically involve direct code analysis, mainly including abstraction interpretation and constraint solving. Abstract interpretation [51] approximates a program’s behavior by mapping its concrete semantics to a simplified abstract domain, often using fixed-point calculations to abstract program states in the presence of loops. Enric et al.[6] developed a framework for generating conjunctions of loop invariants constrained to linear inequalities or polynomial equations, later extending the approach to polynomial invariants of bounded degree[7]. Steven et al. [8] applied a combination of Groebner bases, linear algebra, and abstract interpretation to generate polynomial invariants. Constraint-based methods, on the other hand, rely on pre-defined templates and employ SMT solvers to filter out appropriate invariant expressions. Farkas’ lemma [9, 10, 11, 12, 13] is frequently used to extract linear relationships from loops, while other constraint templates, such as those based on Craig’s interpolant [52, 41, 42], are also employed.
Separation Logic (SL) [36] is an extension of Hoare logic to reason about memory structures and the interactions of programs with dynamically allocated memory. It introduced the concept of "separating conjunction," which allows for concise reasoning about disjoint memory regions and parallel computation. Based on the extension of list manipulation [53], THOR [54, 55, 56] and MemCAD [57] are able to infer loop invariants on numeric and linked list programs. More recently, SLING [38] advances SL invariant generation by employing dynamic analysis to track memory and numerical changes in pointer-related structures.
Learning-based methods Learning-based methods often build upon template-based approaches by learning the parameters within these templates through various machine-learning techniques and ultimately invoking an SMT solver to filter suitable invariants. C2I [58] employs a randomized search to discover candidate invariants, followed by a checker to validate them. ICE [15] did learning for synthesizing invariants, leveraging examples, counterexamples, and implications. Then they combined them with decision trees [59]. CODE2INV [16, 44] utilizes reinforcement learning and graph neural networks to learn program structures. LIPuS [17] also adopts reinforcement learning and uses pruning to reduce the search space. CLN2INV [14] introduces Continuous Logic Networks (CLN) to automatically learn loop invariants directly from program execution traces, while Gated-CLN (G-CLN) [60] extends this approach to allow the model to robustly learn general invariants across a large number of terms.
LLMs for loop invariants Just as we talked about above in the experiment section, LIG-SE [23] fine-tuned the large model and combined it with a checker, giving the LLM the ability to handle separation logic. It can also handle data structures and abstract predicate properties to a certain extent. Autospec [22] employs hierarchical specification generation to produce ACSL-style annotations, which include loop invariants and have achieved good performance in the processing of numerical programs. Both LLM-SE and AutoSpec call the checker at the end to ensure the correctness of the generated invariants. Several other works have explored the intersection of LLMs and loop invariant generation with promising outcomes. The iRank tool [20] introduces a re-ranking approach that evaluates loop invariants produced by LLMs. Unlike other methods, iRank does not focus on generating loop invariants but instead addresses the challenge of verifying the correctness of LLM-generated invariants. To achieve this, the authors employed two distinct LLM-based embedding models to learn the criteria for determining invariant correctness. Kexin et al.[21] aimed to fine-tune a pre-trained LLM for abstract reasoning over program executions, allowing for the generation of loop invariants. Their approach relies on a dynamic analyzer to trace program executions, which are then used to derive invariants. However, their evaluation is limited to a comparison against invariants generated by the dynamic analyzer, which poses inherent constraints. Loopy[24], on the other hand, continuously queries LLMs to generate invariants and employs a verification tool to ensure correctness. While they enhance the generated invariants by cyclically applying the Houdini algorithm and making repairs, the repeated invocation of the verifier introduces significant performance overhead.
@@ 6 Conclusion
Acknowledgment
This work is supported by the National Natural Science Foundation of China Grant No. 61872232.